40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 53 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  14333   Thu Dec 6 17:33:33 2018 JonOmnistructureGeneralN2 line disconnected

I believe I finally have the N2 gauge working correctly. The wiring is unchanged from its original state and the controller has been recalibrated.

After letting the line pressure drop to 0 PSI as indicated by the analog gauge in the drill-press room, I recorded the number of counts read by the Omega controller. Then I pressurized the line to 80 PSI, again indicated by the analog gauge, and recorded the Omega counts again. I entered these two reference points into the controller (automatically determines the gain and offset from these), then confirmed the readings to agree with the anaog gauge as I varied the line pressure.

The two reference points are:

0 PSI  :  13 counts
80 PSI : 972 counts

 

Quote:

[jon, gautam]

In the latest installment in this puzzler: turns out that maybe the trend of the "N2 pressure" channel increasing over the ~3 day timescale it takes a cylinder of N2 to run out is real, and is a feature of the way our two N2 cylinder lines/regulators are setup (for the automatic switching between cylinders when one runs out). In order to test this hypothesis, we'd like to have the line pressure be 0 initially, and then just have 1 cylinder hooked up. 

 

  14332   Thu Dec 6 11:16:28 2018 aaronUpdateOMCOMC channels

I need to hookup +/- 24 V supplies to the OMC whitening/dewhitening boxes that have been added to 1X2.

There are trailing +24V fuse slots, so I will extend that row to leave the same number of slots open.

While removing one +24V wire to add to the daisy chain, I let the wire brush an exposed conductor on the ground side, causing a spark. FSS_PCDRIVE and FSS_FAST are at different levels than before this spark. The 24V sorensens have the same currents as before according to the labels. Gautam advised me to remove the final fuse in the daisy chain before adding additional links.

gautam: we peeled off some outdated labels from the Sorensens in 1X1 such that each unit now has only 1 label visible reflecting the voltage and current. Aaron will post a photo after his work.

  14331   Tue Dec 4 18:24:05 2018 gautamOmnistructureGeneralN2 line disconnected

[jon, gautam]

In the latest installment in this puzzler: turns out that maybe the trend of the "N2 pressure" channel increasing over the ~3 day timescale it takes a cylinder of N2 to run out is real, and is a feature of the way our two N2 cylinder lines/regulators are setup (for the automatic switching between cylinders when one runs out). In order to test this hypothesis, we'd like to have the line pressure be 0 initially, and then just have 1 cylinder hooked up. When we went into the drill-press area, we heard a hiss, turns out that one of the cylinders is leaking (to be fair, this was labelled, but i thought it isn't great to have a higher N2 concentration in an enclosed space). Since we don't need any actuation ability, I valved off the leaky cylinder, and disconnected the other properly functioning one. Attachment #1 shows the current state.

Attachment 1: IMG_7195.JPG
IMG_7195.JPG
  14330   Tue Dec 4 10:38:12 2018 JonOmnistructureUpgradeUpdated Feedthrough List for Vacuum Acromag Chassis

Based on new input from Chub, attached is the revised list of signal cable feedthroughs needed on the vacuum system Acromag crate. I believe this list is now complete.

Attachment 1: acromag_chassis_feedthroughs.pdf
acromag_chassis_feedthroughs.pdf
  14329   Sun Dec 2 19:32:35 2018 ranaUpdateIOOfit times

need to vary start/stop times in fit to test for systematics

  14328   Sun Dec 2 17:26:58 2018 gautamUpdateIMCIMC ringdown fitting

Recently we wondered at the meeting what the IMC round trip loss was. I had done several ringdowns in the winter of 2017, but because the incident light on the cavity wasn't being extinguished completely (the AOM 0th order beam is used), the full Isogaio et. al. analysis could not be applied (there were FSS induced features in the reflection ringdown signal). Nevertheless, I fitted the transmission ringdowns. They looked like clean exponentials, and judging by the reflection signals (see previous elogs in this thread), the first ~20us of data is a clean exponential, so I figured we may get some rough value of the loss by just fitting the transmission data. 

The fitted storage time is 60.8 \pm 2.7 \mu s.However, this number isn't commensurate with the 40m IMC spec of a critically coupled cavity with 2000ppm transmissivity for the input and output couplers.

Attachment #1: Expected storage time for a lossless cavity, with round-trip length ~27m. MC2 is assumed to be perfectly reflecting. The IMC length is known to better than 100 Hz uncertainty because the marconi RF modulation signal is set accordingly. For the 40m spec, I would expect storage times of ~40 usec, but I measure almost 30% longer, at ~60 usec.

Attachment #2: Fits and residuals from the 10 datasets I had collected. This isn't a super informative plot because there are 10 datasets and fits, but to eye, the fits are good, and the diagonal elements of the covariance matrix output by scipy's curve_fit back this up. The function used to fit the t > 0 portions of these signals (because the light was extinguished at t=0 by actuating on the AOM) is \text{Transmission} = Ae^{-\frac{2t}{\tau_{\mathrm{storage}}}}, where A and tau are the fitted parameters. In the residuals, the same artefacts visible in the reflection signal are seen.

Attachment #3: Scatter plot of the data. Width of circles are proportional to fit error on individual measurements (i just scaled the marker size arbitrarily to be able to visually see the difference in uncertainty, the width doesn't exactly indicate the error), while the dahsed lines are the global mean and +/- 1 sigma levels.

Attachment #4: Cavity pole measurement. Using this, I get an estimate of the loss that is a much more believable 300 \pm 20\, \mathrm{ppm}.

Attachment 1: tauTheoretical.pdf
tauTheoretical.pdf
Attachment 2: ringdownFit.pdf
ringdownFit.pdf
Attachment 3: ringdownScatter.pdf
ringdownScatter.pdf
Attachment 4: cavPole.pdf
cavPole.pdf
  14327   Sun Dec 2 16:08:44 2018 JonOmnistructureUpgradeFeedthroughs for Vacuum Acromag Chassis

Below is an inventory of the signal feedthroughs that need to be installed on the vacuum Acromag crate this week.

Type Qty Connects to # Chs Signals
DB-37 female 1 Main AC relay box 18 Valve/roughing pump control
DB-9 female 5** Satellite AC relay boxes 3-4/box Valve control
DB-25 male 1 Turbo pump 1 controller 5 Pump status readbacks
DB-9 male 30 Valve position indicators 2/valve Valve position readbacks
DB-9 male 3 Roughing pump controllers 1/pump Pump status readbacks
DB-9 male 1 Cryo pump controller 2 Pump status readbacks

**The original documentation lists five satellite boxes (one for each test mass chamber and one for the beamsplitter chamber), but Chub reports not all of them are in use. We may remove the ones not used.

  14326   Fri Nov 30 19:37:47 2018 gautamUpdateLSCLSC feedforward block diagram

I wanted to set up an RTCDS model to understand this problem better. Attachment #1 is the simulink diagram of the signal flow. The idea will be to put in the appropriate filter shapes into the various filter blocks denoting the DARM and auxiliary DoF plants, controllers and actuators, and then use awggui / diaggui to inject some noises and see if in this idealized model I can achieve good subtraction. Then we can build up to applying a time varying cross coupling between DARM and the vertex DoF, and see how good the adaptive FF works. Still need to setup some MEDM screens to make working with the test system easier.

I figured c1omc would be the least invasive model to set this upon without risking losing any of our IR/green alignment references. Compile and install went smooth, see Attachment #2. The c1omc model was clocking 4us before, now it's using 7us.

Attachment #3 shows the top level of the OMC model, while Attachment #4 shows the MEDM screen.

* Note to self: when closing a loop inside the realtime model, there has to be a delay block somewhere in the loop, else a compilation error is thrown.

Attachment 1: LSC_FF_tester.png
LSC_FF_tester.png
Attachment 2: Screenshot_from_2018-11-30_19-41-07.png
Screenshot_from_2018-11-30_19-41-07.png
Attachment 3: Screenshot_from_2018-12-10_15-31-23.png
Screenshot_from_2018-12-10_15-31-23.png
Attachment 4: SimLSC.png
SimLSC.png
  14325   Fri Nov 30 15:53:52 2018 JonOmnistructureGeneralN2 pressure gauge fix

I've made a repair to the N2 pressure monitor. I don't believe the polarity of the analog signal into the controller actually was reversed. I found the data sheet (attached) for the transducer model we have installed. Its voltage should read ~0 at 0 PSI and 100mV at 100 PSI. As wired, the input voltage reads +80 mV as it should.

The controller calibrates the sensor voltage to PSI (i.e., applies a scale and offset) based on two settable reference points which appeared to be incorrect. I changed them to:

  1. 0 mV = 0 PSI (neglecting the small dark bias)
  2. 100 mV = 100 PSI

After the change, the pressure reads 80 PSI. Let's see if the time history now shows a sensible trend. 

Quote:

[koji, gautam, jon, steve]

  • We suspect analog voltage from N2 pressure gauge is connected to interfacing Omega controller with the 'wrong' polarity (i.e pressure is rising over ~4 days and then rapidly falling instead of the other way around). This should be fixed.

 

Attachment 1: N2_transducer_datasheet.pdf
N2_transducer_datasheet.pdf
  14324   Thu Nov 29 17:46:43 2018 gautamUpdateGeneralSome to-dos

[koji, gautam, jon, steve]

  • We suspect analog voltage from N2 pressure gauge is connected to interfacing Omega controller with the 'wrong' polarity (i.e pressure is rising over ~4 days and then rapidly falling instead of the other way around). This should be fixed.
  • N2 check script logic doesn't seem robust. Indeed, it has not been sending out warning emails (threshold is set to 60 psi, it has certainly gone below this threshold even with the "wrong" polarity pressure gauge hookup). Probably the 40m list is rejecting the email because controls isn't a part of the 40m group.
  • Old frames have to be re-integrated from JETSTOR to the new FB in order to have long timescale lookback.
  • N2 cylinder pressure gauges (at the cylinder end) need a power supply - @ Steve, has this been purchased? If not, perhaps @ Chub can order it.
  • MEDM vacuum screen should be updated to have gate valves be a different color to the spring-loaded valves. Manual valve between TP1 and V1 should also be added.
  • P2, P3 and P4 aren't returning sensible values (they should all be reading ~760torr as is P1). @ Steve, any idea if these gauges are broken?
  • Hornet gauges (CC and Pirani) should be hooked up to the new vacuum system.
  • add slow channels of   foreline pressures of TP2 & 3   and    C1:Vac-IG1_status_pressure
  14323   Thu Nov 29 08:13:33 2018 SteveUpdatePEMEQ 3.9m So CA

EQ did not trip anything. atm1

Just a REMINDER: our vacuum system is at atm to help the vacuum upgrade to Acromag.

Exceptions: cryo pump and 4 ion pumps

It is our first rainy day of the season..The roof is not leaking.

 Vac Status: The vac rack power was recycled yesterday and power to controller TP1,2 and 3 restored. atm3

                     VME is OFF.        Power to all other instrument are ON.       23.9Vdc 0.2A

ETMY sus tower with locked optic in HEPA tent at east end is standing by for action.

 

Attachment 1: 3.9mSoCA.png
3.9mSoCA.png
Attachment 2: Vac_as_today.png
Vac_as_today.png
Attachment 3: as_is.png
as_is.png
  14322   Tue Nov 27 17:06:51 2018 SteveConfigurationVACAgilent 84FS turbo installed as TP2

Chub & Steve,

We swapped in our  replacement of Varian V70D "bear-can" turbo as factory clean.

The new Agilent TwisTorr 84 FS  turbo pump [ model x3502-64002,  sn IT17346059 ]  with intake screen, fan, vent valve. The controller  [ model 3508-64001, sn IT1737C383 ] and a larger drypump IDP-7,  [ model x3807-64010, sn MY17170019 ] was installed.

Next things to do:

  1. implement hardware interlock to close V4 at 80% pumping speed slowdown of "standby" rotation speed, estimated to be ~ 40,000 RPM ( when Standby 50K RPM  )
  2. set up isolation valve in the foreline of TP2, with delayed start of the IDP-7 and/or use relay to power drypump.  This turbo controller can not switch off or start of the dry pump. [ Agilent isolation valve #X3202-60055, with position indicator, pneumatic actuation, 115V solenoid ]..........as a second thought, we do not need isolation valve if we go with the relay option. The IDP-7 has built in delay of 10-15 sec
  3. test performance of new turbo
  14321   Tue Nov 27 10:50:20 2018 SteveUpdatePEMearth quake Mexico

Nothing tripped.

 

Attachment 1: 5.5M.Mexico.png
5.5M.Mexico.png
  14320   Mon Nov 26 21:58:08 2018 JonOmnistructure Serial Vacuum Signals

All the serial vacuum signals are now interfaced to the new digital controls system. A set of persistent Python scripts will query each device at regular intervals (up to ~10 Hz) and push the readings to soft channels hosted by the modbus IOC. Similar scripts will push on/off state commands to the serial turbo pumps.

IP Addresses/Comm Settings

Each serial device is assigned an IP address on the local subnet as follows. Its serial communication parameters as configured in the terminal server are also listed.

Device IP Address Baud Rate Data Bits Stop Bits Parity
MKS937a vacuum gauge controller 192.168.114.11 9600 8 1 even
MKS937b vacuum gauge controller 192.168.114.12 9600 8 1 even
GP307 vacuum gauge controller 192.168.114.13 9600 8 1 even
GP316a vacuum gauge controller 192.168.114.14 9600 8 1 even
GP316b vacuum gauge controller 192.168.114.15 9600 8 1 even
N2 pressure line gauge 192.168.114.16 9600 7 1 odd
TP2/3 192.168.114.17/18 9600 8 1 none

Hardware Modifications

  • Each of the five vacuum gauge controllers has an RJ45 adapter installed directly on its DB9/DB25 output port. Because the RJ45 cable now plugs directly into the terminal server, instead of passing through some additional adapters as it formerly did, it was necessary to reverse the wiring of the controller TXD and RXD pins to Ethernet pins. The DB9/25-to-RJ45 adapters on the back of the controllers are now wired as follows.
    • For the MKS controllers: DB2 (RXD) --> Eth4;  DB3 (TXD) --> Eth5;  DB5 (RTN) --> Eth6
    • For the Granville-Phillips controllers: DB2 (TXD) --> Eth5;  DB3 (RXD) --> Eth4;  DB7 (RTN) --> Eth6
  • I traced a communications error with the GP307 gauge controller all the way back to what I would have suspected least, the controller itself. The comm card inside each controller has a set of mechanical relay switches which set the communications parameters (baud rate, parity, etc.). Knowing that this controller was not part of the original installation, but was swapped in to replace the original in 2009, I pulled the controller from the rack and checked the internal switch settings. Sure enough, the switch settings (pictured below) were wrong. In the background of the photo is the unit removed in 2009, which has the correct settings. After setting the correct communications parameters, the controller immediately began communicating with the server. Did these readouts (PRP, PTP1) never work since 2009? I don't see how they could.
Attachment 1: GP307_relays.jpeg
GP307_relays.jpeg
  14319   Mon Nov 26 17:16:27 2018 gautamUpdateSUSEY chamber work

[steve, rana, gautam]

  • PSL and EY 1064nm laser (physical) shutters on the head were closed so that we and sundance crew could work without laser safety goggles. EY oplev laser was also turned off.
  • Cylindrical heater setup removed:
    • heater wiring meant the heater itself couldn't be easily removed from the chamber
    • two lenses and Al foil cylinder removed from chamber, now placed on the mini-cleanroom table.
  • Parabolic heater is untouched for now. We can re-insert it once the test mass is back in, so that we can be better informed about the clipping situation.
  • ETMY removed from chamber.
    • EQ stops were engaged.
    • Pictures were taken
    • OSEMs were removed from cage, placed in foil holders.
    • Cage clamps were removed after checking that marker clamps were in place.
    • Optic was moved first to NW corner of table, then out of the vacuum onto the mini-cleanroom desk Chub and I had setup last week.
    • Hoepfully there isn't an earthquake. EY has been marked as off-limits to avoid accidental bumping / catasrophic wire/magnet/optic breaking.
    • We sealed up the mini cleanroom with tape. F.C. cleaning tomorrow or at another opportune moment.
    • Light door was put back on for the evening.

Rana pointed out that the OSEM cabling, because of lack of a plastic shielding, is grounded directly to the table on which it is resting. A glass baking dish at the base of the seismic stack prevents electrical shorting to the chamber. However, there are some LEMO/BNC cables as well on the east side of the stack, whose BNC ends are just lying on the base of the stack. We should use this opportunity to think about whether anything needs to be done / what the influence of this kind of grounding is (if any) on actuator noise.

Steve also pointed out that we should replace the rubber pads which the vacuum chamber is resting on (Attachment #1, not from this vent, but just to indicate what's what). These serve the purpose of relieving small amounts of strain the chamber may experience relative to the beam tube, thus helping preserve the vacuum joints b/w chamber and tube. But after (~20?) years of being under compression, Steve thinks that the rubber no longer has any elasticity, and so should be replaced.

Attachment 1: IMG_5251.JPG
IMG_5251.JPG
  14318   Mon Nov 26 15:58:48 2018 SteveUpdateVACVent 81

Gautam, Aaron, Chub & Steve,

ETMY heavy door replaced by light one.

We did the following:  measured 950 particles/cf min of 0.5 micron at SP table, wiped crane and it's cable, wiped chamber,

                                placed heavy door on clean merostate covered stand, dry wiped o-rings and isopropanol wiped Aluminum light cover

                              

Quote:

Gautam, Aaron, Chub and Steve,

Quote:

Vent 80 is nearly complete; the instrument is almost to atmosphere.  All four ion pump gate valves have been disconnected, though the position sensors are still connected,and all annulus valves are open.  The controllers of TP1 and TP3 have been disconnected from AC power. VC1 and VC2 have been disconnected and must remained closed. Currently, the RGA is being vented through the needle valve and the RGA had been shut off at the beginning of the vent preparations.  VM1 and VM3 could not be actuated.  The condition status is still listed as Unidentified because of the disconnected valves. 

The vent 81 is completed.

4 ion pumps and cryo pump are at ~ 1-4 Torr (estimated as we have no gauges there), all other parts of the vacuum envelope are at atm. P2 & P3 gauges are out of order.

V1 and VM1 are in a locked state. We suspect this is because of some interlock logic.

TP1 and TP3 controllers are turned off.

Valve conditions as  shown: ready to be opened or closed or moved or rewired. To re-iterate: VC1, VC2, and the Ion Pump valves shouldn't be re-connected during the vac upgrade.

Thanks for all of your help.

 

  14317   Mon Nov 26 15:43:16 2018 aaronUpdateOMCOMC scanning/aligning script

I've started testing the OMC channels I'll use.

I needed to update the model, because I was getting "Unable to setup testpoint" errors for the DAC channels that I had created earlier, and didn't have any ADC channels yet defined. I attach a screenshot of the new model. I ran

rtcds make c1omc
rtcds install c1omc
rtcds start c1omc.
 
without errors.
Attachment 1: c1omc.png
c1omc.png
  14316   Mon Nov 26 10:22:16 2018 aaronUpdateGeneralprojector light bulb replaced

I replaced the projector bulb. Previous bulb was shattered.

  14315   Sun Nov 25 17:41:43 2018 JonOmnistructure Vacuum Controls Upgrade - Status and Plans

New hardware has been installed in the vacuum controls rack. It is shown in the below post-install photo.

  • Supermicro server (c1vac) which will be replacing c1vac1 and c1vac2.
  • 16-port Ethernet switch providing a closed local network for all vacuum devices.
  • 16-port IOLAN terminal server for multiplexing/Ethernetizing all RS-232 serial devices.

Below is a high-level summary of where things stand, and what remains to be done.

Completed:

 Set up of replacement controls server (c1vac).

  • Supermicro 1U rackmount server, running Debian 8.5.
  • Hosting an EPICS modbus IOC, scripted to start/restart automatically as a system service.
  • First Ethernet interface put on the martian network at 192.168.113.72.
  • Second Ethernet interface configured to host a LAN at 192.168.114.xxx for communications with all vacuum electronics. It connects to a 16-port Ethernet switch installed in the vacuum electronics rack.
  • Server installed in vacuum electronics rack (see photo).

 Set up of Acromag terminals.

  • 6U rackmount chassis frame assembled; 15V DC, 24V DC, and Ethernet wired.
  • Acromags installed in chassis and configured for the LAN (5 XT1111 units, 2 XT1121 units).

 EPICS database migration.

  • All vacuum channels moved to the modbus IOC, with the database updated to address the new Acromags. [The new channels are running concurrently at "C1:Vac2-...." to avoid conflict with the existing system.]
  • Each hard channel was individually tested on the electronics bench to confirm correct addressing and Acromag operation.

 Set up of 16-port IOLAN terminal server (for multiplexing/Ethernetizing the serial devices).

  • Configured for operation on the LAN. Each serial device port is assigned a unique IP address, making the terminal server transparent to client TCP applications.
  • Most of the pressure gauges are now communicating with the controls server via TCP.

Ongoing this week:

  • [Jon] Continue migrating serial devices to ports on the terminal server. Still left are the turbo pumps, N2 gauge, and RGA.
  • [Jon] Continue developing Python code for communicating with gauges and pumps via TCP sockets. A beta version of gauge readout code is running now.
  • [Chub] Install feedthrough panels on the Acromag chassis. Connect the wiring from feedthrough panels to the assigned Acromag slots.
  • [Chub/Jon] Test all the hard EPICS channels on the electronics bench, prior to installing the crate in the vacuum rack.
  • [Chub/Jon] Install the crate in the vacuum rack; connect valve/pump readbacks and actuators; test each hard EPICS channel in situ.
  • [Jon] Once all the signal connections have been made, in situ testing of the Python interlock code can begin.
Attachment 1: rack_photo.jpg
rack_photo.jpg
  14314   Wed Nov 21 16:48:11 2018 gautamUpdateCOCEY mini cleanroom setup

With Chub's help, I've setup a mini cleanroom at EY - Attachment #1. The HEPA unit is running on high now. All surfaces were wiped with isopropanol, we can wipe everything down again on Monday and replace the foil.

Attachment 1: IMG_7174.JPG
IMG_7174.JPG
  14313   Wed Nov 21 09:59:26 2018 gautamUpdateLSCLSC feedforward block diagram

Attachment #1 is a block diagram depicting the pathway by which the vertex DOF control signals can couple into DARM (adapted from a similar diagram in Gabriele's Virgo note on the subject). I've also indicated some points where noise can couple into either loop. In general, there are sensing noises that couple in at the error point of the loop, and actuation noises that couple in at the control point. In this linear picture, each block represents a (possibly time varying) transfer function. So we can write out the node-to-node transfer functions and evaluate the various couplings.

The motivation is to see if we can first simulate with some realistic noise and time-varying couplings (and then possibly test on the realtime system) the effectiveness of the filter denoted by "FF" in canceling out the shot noise from the auxiliary loop being re-injected into the DARM loop via the DARM sensor. Does this look correct?

Attachment 1: IMG_7173.JPG
IMG_7173.JPG
  14312   Tue Nov 20 20:33:11 2018 aaronUpdateOMCOMC scanning/aligning script

I finished running the cabling for the OMC, which involved running 7x 50ft DB9 cables from the OMC_NORTH rack to the 1X2 rack, laying cables over others on the tray. I tried not to move other cables to the extent I could, and I didn't run the new cables under any old cables. I attach a sketch diagram of where these cables are going, not inclusive of the entire DAC/ADC signal path.

I also had to open up the AA board (D050387, D050374), because it had an IPC connector rather than the DB37 that I needed to connect. The DAC sends signals to a breakout board that is in use (D080302) and had a DB37 output free (though note this carries only 4 DAC channels). I opened up the AA board and it had two IPC 40s connected to an adapter to the final IPC 70 output. I replaced the IPC40 connectors with DB37 breakouts, and made a new slot (I couldn't find a DB37 punch, so this is not great...) on the front panel for one of them, so I can attach it to the breakout board.

I noticed there were many unused wires, so I had to confirm that I had the wiring correct (still haven't confirmed by driving the channels, but will do). There was no DCC for D080302, but I grabbed the diagrams for the whitening boards it was connected to (D020432) and for the AA board I was opening up as well as checked out elog 8814, and I think I got it. I'll confirm this manually and make a diagram if it's not fake news.

Attachment 1: pathwaysketch.pdf
pathwaysketch.pdf
Attachment 2: IMG_0094.JPG
IMG_0094.JPG
Attachment 3: IMG_0097.JPG
IMG_0097.JPG
  14311   Tue Nov 20 17:38:13 2018 ranaUpdateUpgradeNew Coffee Machine

Rana, Aaron, Gautam

The old Zojirushi has died. We have received and comissioned our new Technivoorm Mocha Master today. It is good.

  14310   Tue Nov 20 13:13:01 2018 gautamUpdateVACIMC alignment is okay

I checked the IMC alignment following the vent, for which the manual beam block placed on the PSL table was removed. The alignment is okay, after minor touchup, the MC Trans was ~1200 cts which is roughly what it was pre-vent. I've closed the PSL shutter again.

  14309   Mon Nov 19 23:38:41 2018 JonOmnistructure Vacuum Acromag Channel Assignments

I've completed bench testing of all seven vacuum Acromags installed in a custom rackmount chassis. The system contains five XT1111 modules (sinking digital I/O) used for readbacks of the state of the valves, TP1, CP1, and the RPs. It also contains two XT1121 modules (sourcing digital I/O) used to pass 24V DC control signals to the AC relays actuating the valves and RPs. The list of Acromag channel assignments is attached.

I tested each input channel using a manual flip-switch wired between signal pin and return, verifying the EPICS channel readout to change appropriately when the switch is flipped open vs. closed. I tested each output channel using a voltmeter placed between signal pin and return, toggling the EPICS channel on/off state and verifying the output voltage to change appropriately. These tests confirm the Acromag units all work, and that all the EPICS channels are correctly addressed.

Attachment 1: Binary_IO_Channel_Assignments.pdf
Binary_IO_Channel_Assignments.pdf
  14308   Mon Nov 19 22:45:23 2018 JonOmnistructure Vacuum System Subnetwork

I've set up a closed subnetwork for interfacing the vacuum hardware (Acromags and serial devices) with the new controls machine (c1vac; 192.168.113.72). The controls machine has two Ethernet interfaces, one which faces outward into the martian network and another which faces the internal subnetwork, 192.168.114.xxx. The second network interface was configured via the following procedure.

1. Add the following lines to /etc/network/interfaces:

allow-hotplug eth1
iface eth1 inet static
address 192.168.114.9
netmask 255.255.255.0

2. Restart the networking services:

$sudo /etc/init.d/networking restart

3. Enable DNS lookup on the martian network by adding the following lines to /etc/resolv.conf:

search martian
nameserver 192.168.113.104

4. Enable IP forwarding from eth1 to eth0:

$sudo echo 1 > /proc/sys/net/ipv4/ip_forward

5. Configure IP tables to allow outgoing connections, while keeping the LAN invisible from outside the gateway (c1vac):

$sudo iptables -t nat -A POSTROUTING -o eth0 -j MASQUERADE
$sudo iptables -A FORWARD -i eth0 -o eth1 -m state --state RELATED,ESTABLISHED -j ACCEPT
$sudo iptables -A FORWARD -i eth1 -o eth0 -j ACCEPT

6. Finally, because the EPICS 3.14 server binds to all network interfaces, client applications running on c1vac now see two instances of the EPICS server---one at the outward-facing address and one at the LAN address. To resolve this ambiguity, two additional enviroment variables must be set that specify to local clients which server address to use. Add the following lines to /home/controls/.bashrc:

EPICS_CA_AUTO_ADDR_LIST=NO
EPICS_CA_ADDR_LIST=192.168.113.72

A list of IP addresses so far assigned on the subnetwork follows.

Device IP Address
Acromag XT1111a 192.168.114.1
Acromag XT1111b 192.168.114.2
Acromag XT1111c 192.168.114.3
Acromag XT1111d 192.168.114.4
Acromag XT1111e 192.168.114.5
Acromag XT1121a 192.168.114.6
Acromag XT1121b 192.168.114.7
Perle IOLAN SDS16 192.168.114.8
c1vac 192.168.114.9
  14307   Mon Nov 19 22:01:50 2018 gautamUpdateVACLoose nut on valve

As I was turning off the lights in the VEA, I heard a rattling sound from near the PSL enclosure. I followed it to a valve - I couldn't see a label on this valve in my brief effort to find one, but it is on the south-west corner of the IMC table, so maybe VABSSCI or VABSSCO? The power cable is somehow spliced with an attachment that looks to be bringing gas in/out of the valve (See Attachment #1), and the nut on the bottom was loose, the whole power cable + mettal attachment was responsible for the rattling. I finger-tightened the nut and the sound went away.

Attachment 1: IMG_7171.JPG
IMG_7171.JPG
  14306   Mon Nov 19 17:09:00 2018 SteveUpdateVACVent 81

Gautam, Aaron, Chub and Steve,

Quote:

Vent 80 is nearly complete; the instrument is almost to atmosphere.  All four ion pump gate valves have been disconnected, though the position sensors are still connected,and all annulus valves are open.  The controllers of TP1 and TP3 have been disconnected from AC power. VC1 and VC2 have been disconnected and must remained closed. Currently, the RGA is being vented through the needle valve and the RGA had been shut off at the beginning of the vent preparations.  VM1 and VM3 could not be actuated.  The condition status is still listed as Unidentified because of the disconnected valves. 

The vent 81 is completed.

4 ion pumps and cryo pump are at ~ 1-4 Torr (estimated as we have no gauges there), all other parts of the vacuum envelope are at atm. P2 & P3 gauges are out of order.

V1 and VM1 are in a locked state. We suspect this is because of some interlock logic.

TP1 and TP3 controllers are turned off.

Valve conditions as  shown: ready to be opened or closed or moved or rewired. To re-iterate: VC1, VC2, and the Ion Pump valves shouldn't be re-connected during the vac upgrade.

Thanks for all of your help.

Attachment 1: beforeVent82.png
beforeVent82.png
Attachment 2: vent81completed.png
vent81completed.png
  14305   Mon Nov 19 14:59:48 2018 ChubUpdateVACVent 81

Vent 80 is nearly complete; the instrument is almost to atmosphere.  All four ion pump gate valves have been disconnected, though the position sensors are still connected,and all annulus valves are open.  The controllers of TP1 and TP3 have been disconnected from AC power. VC1 and VC2 have been disconnected and must remained closed. Currently, the RGA is being vented through the needle valve and the RGA had been shut off at the beginning of the vent preparations.  VM1 and VM3 could not be actuated.  The condition status is still listed as Unidentified because of the disconnected valves. 

  14304   Sun Nov 18 17:09:02 2018 gautamUpdateGeneralVent prep

Vent prep

Following the checklist, I did these:

  • Both arms were locked to IR, TRY and TRX maximized using ASS.
  • GTRY and GTRX were also maximized.
  • ITM/ETM Oplevs centered with TRX/TRY maximized, PRM/SRM/BS Opelvs were centered once the DRMI was locked and aligned.
  • Attachment #1 summarizes the above 3 bullets.
  • Sensoray was made to work with Donatella (Raspberry Pi video server idea is good but although the sensoray drivers look to have installed correctly, when the Sensoray unit is plugged into the RPi USB port, the red light doesn't come on, so I opted to not spend too much time on it for the moment).
  • Photos of all ports in various locked configurations are saved in /users/sensoray/SensorayCaptures/Nov2018Vent
  • PSL power into the IMC was cut from 1.07 W (measured after G&H mirror) to 97 mW. I opted to install a new HWP+PBS after the PMC to cut the power, so we don't have to fiddle around so much with the PMC locking settings [Attachment #3, this was the only real candidate location as the IMC wants s-polarized light].
  • 2" R=10% BS in the IMC REFL path was replaced with a 2" Y1 HR mirror, so there is no MCREFL till we turn the power back up.
  • IMC was locked.
  • Low power MC autolocker works [Attachment #2]. The reduction in MCREFL is because of me manually aligning the cavity, WFS servos are disabled in low power mode since there is no light incident on the WFS heads.
  • Updated the SUS driftmon values (though I'm not really sure how useful this is).
  • PSL shutter will remain closed, but I have not yet installed a manual beam block in the beam path on the PSL table.

@Steve & Chub, we are ready to vent tomorrow (Monday Nov 19). 

Attachment 1: VentPrepNov2018.png
VentPrepNov2018.png
Attachment 2: MCautolocker_lowPower.png
MCautolocker_lowPower.png
Attachment 3: IMG_7163.JPG
IMG_7163.JPG
  14303   Sun Nov 18 00:59:33 2018 gautamUpdateGeneralVent prep

I've begun prepping the IFO for the vent, and completed most of the IFO related items on the checklist. The power into the MC has been cut, but the low-power autolocker has not been checked. I will finish up tomorrow and post the go ahead. PSL shutter is closed for tonight.

  14302   Sat Nov 17 18:59:01 2018 aaronUpdateIOOIMC problematic

I made additional measurements on the x and y arms, at 5 offset positions for each arm (along with 6 measurements at the "zeroed" position).

  14301   Fri Nov 16 15:09:31 2018 SteveConfigurationVACnot venting cryo and ion pumps

Notes on the ion pumps and cryo pump:

  • Our 4 ion pumps were closed off for a lomg time. I estmated their pressure to be around ~1 Torr. After talking with Koji we decided not to vent them.

  • It'd be still useful to wire their position sensors. But make sure we do not actuate the valves. 

  • The cryo pump was regenerated to 1e-4 Torr about 2 years ago. It's pressure can be ~ 2 Torr with charcoal powder. It is a dirty system at room temperature.

  • Do not actuate VC1 and VC2, and keep its manual valve closed.

  • IF someone feels we should vent them for some reason, let us know here in the elog before Monday morning.

 

Quote:

Wiring of the power, Ethernet, and indicator lights for the vacuum Acromag chassis is complete. Even though this crate will only use +24V DC, I wired the +/-15V connector and indicator lights as well to conform to the LIGO standard. There was no wiring diagram available, so I had to reverse-engineer the wiring from the partially complete c1susaux crate. Attached is a diagram for future use. The crate is ready to begin software developing on Monday. 

 

  14300   Fri Nov 16 10:53:07 2018 aaronUpdateIOOIMC problematic

Back to loss measurements.

I replaced the PD I've been using for the AS beam.

I misaligned the x arm.

I tried to lock the y arm, but PRC was locked so I could was unable. Gautam reminded me where the config scripts are.

The armloss measurement script needed two additional modifications:

  • It was setting the initial offset of the PIT and YAW demod signals to 0, but due to the clipping on the heater we are operating at an offset. I commented out these lines.
  • When the script ran UNFREEZE_DITHER, it was running it using medmrun. The scope script hadn't been using this, and it seemed that when it ran UNFREEZE_DITHER in this way the YARM_ASS servo was passing only '0'. I don't really know why this was, but when I removed the call to medmrun it worked.

I ran successfully the loss measurement script for the x and y arms. I'm getting losses of ~100ppm from the first estimates.

I made the following changes to the lossmap script:

  • make the averaging time an input to the script, so we can exceed 2 second averages
  • remove anything about getting data from the scope, replace it with the correct analogues to save the averages for POX/POY refl, MC trans, op lev P/Y, and ASDC signal.
  • record the GPS time in the file with the cds averages (this way I can grab the full data)
  • Added a step in the lossmap script to misalign the optic, so we can continue getting data for the 'misaligned' state, both for the centered and not-centered measurements (that is, for every position on the lossmap).

When the optic aligns itself not at the ideal position, I'm noticing that it often locks on a 01. When the cavity is then misaligned and restored, it can no longer obtain lock. To fix this, I've moved my 'save' commands to just before the loop begins. This means the script may take longer to run, but as long as the cavity is initially locked and well aligned, this should make it more robust against wandering off and never reacquiring lock.

I left the lossmap script running for the x-arm. Next would be to run it for the y arm, but I see that after stepping to a few positions the lock is again lost. It's still trying to run, but if you want to stop it no data already taken will be lost. To stop it, go to the remaining terminal open on rossa and ctrl+c

the analysis needs:

  • Windowing
  • Filter, don't average
  • detrend to get rid of the linear drifts in lock that we see.
    • Is this the right thing?
Attachment 1: Screenshot_from_2018-11-16_19-22-34.png
Screenshot_from_2018-11-16_19-22-34.png
  14299   Fri Nov 16 10:26:12 2018 SteveUpdateVAC single viton O-rings

The 40m vacuum envelope has one large single O-ring on the OOC west side. All other doors have double O-ring with annuloses.

There are 3  spacers to protect o-ring. They should not be removed!

 

The Cryo-pump static seal to VC1 also viton. All gate valves and right angle valve plates have single viton o-ring seal.

Small single viton o-rings on all optical quality viewports.

Helium will permiate through these fast. Leak checking time is limited to 5-10 minutes.

All other seals are copper gaskits. We have 2 manual right angle with METAL-dynamic seal [ VATRING ] as  VV1 & RV1

 

Attachment 1: Single-O-ring.png
Single-O-ring.png
  14298   Fri Nov 16 00:47:43 2018 gautamUpdateLSCMore DRMI characterization

Summary:

  • More DRMI characterization was done.
  • I was working on trying to improve the stability of the DRMI locks as this is necessary for any serious characterization.
  • Today I revived the PRC angular feedforward - this was a gamechanger, the DRMI locks were much more stable. It's probably worth spending some time improving the POP LSC/ASC sensing optics/electronics looking towards the full IFO locking.
  • Quantitatively, the angular fluctuations as witnessed by the POP QPD is lowered by ~2x with the FF on compared to offyes [Attachment #1, references are with FF off, live traces are with FF on].
  • The first DRMI lock I got is already running 15 mins, looking stable.
    • Update: Out of the ~1 hour i've tried DRMI locking tonight, >50 mins locked!
  • I think the filters can be retrained and this performance improved, something to work on while we are vented.
  • Ran sensing lines, measured loop TFs, analysis tomorrow, but I think the phasing of the 1f PDs is now okay.
    • MICH in AS55 Q, demod phase = -92deg, +6dB wht gain.
    • PRCL in REFL11 I, demod phase = +18 deg, +18dB wht gain.
    • SRCL in REFL55 I, demod phase = -175 deg, +18dB wht gain.
  • Also repeated the line in SRCL-->witness in MICH test.
    • At least 10 minutes of data available, but I'm still collecting since the lock is holding.
    • This time I drove the line at ~124 Hz with awggui, since this is more a regime where we are sensing noise dominated.

Prep for this work:

  • Reboots of c1psl, c1iool0, c1susaux.
  • Removed AS port PD loss measurement PD.
  • Initial alignment procedure as usual: single arms --> PRMI locked on carrier --> DRMI

I was trying to get some pics of the optics as a zeroth-level reference for the pre-vent loss with the single arms locked, but since our SL7 upgrade, the sensoray won't work anymore no. I'll try fixing this during the daytime.

Attachment 1: PRCff.pdf
PRCff.pdf
Attachment 2: DRMI_withPRCff.png
DRMI_withPRCff.png
  14297   Thu Nov 15 10:21:07 2018 aaronUpdateIOOIMC problematic

I ran a BNC from the PD on the AS table along the cable rack to a free ADC channel on the LSC whitening board. I lay the BNC on top of the other cables in the rack, so as not to disturb anything. I also was careful not to touch the other cables on the LSC whitening board when I plugged in my BNC. The PD now reads out to... a mystery channel. The mystery channel goes then to c1lsc ADC0 channels 9-16 (since the BNC goes to input 8, it should be #16). To find the channel, I opened the c1lsc model and found that adc0 channel 15 (0-indexed in the model) goes to a terminator.

Rather than mess with the LSC model, Gautam freed up C1:ALS-BEATY_FINE_I, and I'm reading out the AS signal there.

I misaligned the x-arm then re-installed the AS PO PD, using the scope to center the beam then connecting it to the BNC to (first the mystery channel, then BEATY). I turned off all the lights.

I went to misalign the x-arms, but the some of the control channels are white boxed. The only working screen is on pianosa.

The noise on the AS signal is much larger than that on the MC trans signal, and the DC difference for misaligned vs locked states is much less than the RMS (spectrum attached); the coherence between MC trans and AS is low. However, after estimating that for ~30ppm the locked vs misaligned states should only be ~0.3-0.4% different, and double checking that we are well above ADC and dark noise (blocked the beam, took another spectrum) and not saturating the PD, these observations started to make more sense.

To make the measurement in cds, I also made the following changes to a copy opf Johannes' assess_armloss_refl.py that I placed in /opt/rtcds/caltech/c1/scripts/lossmap_scripts/armloss_cds/   :

  • function now takes as argument the number of averages, averaging time, channel of the AS PD, and YARM|XARM|DARK.
  • made the data save to my directory, in /users/aaron/40m/data/armloss/

I started taking a measurement, but quickly realized that the mode cleaner has been locked to a higher order mode for about an hour, so I spend some time moving the MC. It would repeatedly lock on the 00 mode, but the alignment must be bad because the transmission fluctuates between 300 and 1400, and the lock only lasts about 5 minutes.

Attachment 1: 181115_chansDown.png
181115_chansDown.png
Attachment 2: PD_noise.png
PD_noise.png
  14296   Wed Nov 14 21:34:44 2018 JonOmnistructure Vacuum Acromags installed and tested

All 7 Acromag units are now installed in the vacuum chassis. They are connected to 24V DC power and Ethernet.

I have merged and migrated the two EPICS databases from c1vac1 and c1vac2 onto the new machine, with appropriate modifications to address the Acromags rather than VME crate.

I have tested all the digital output channels with a voltmeter, and some of the inputs. Still more channels to be tested.

I’ll follow up with a wiring diagram for channel assignments.

Attachment 1: IMG_3003.jpg
IMG_3003.jpg
  14295   Wed Nov 14 18:58:35 2018 aaronUpdateDAQNew DAC for the OMC

I began moving the AA and AI chassis over to 1X1/1X2 as outlined in the elog.

The chassis were mostly filled with empty cables. There was one cable attached to the output of a QPD interface board, but there was nothing attached to the input so it was clearly not in use and I disconnected it.

I also attach a picture of some of the SMA connectors I had to rotate to accommodate the chassis in their new locations.

Update:

The chassis are installed, and the anti-imaging chassis can be seen second from the top; the anti-aliasing chassis can be seen 7th from the top.

I need to breakout the SCSI on the back of the AA chassis, because ADC breakout board only has a DB36 adapter available; the other cables are occupied by the signals from the WFS dewhitening outputs.

Attachment 1: 6D079592-1350-4099-864B-1F61539A623F.jpeg
6D079592-1350-4099-864B-1F61539A623F.jpeg
Attachment 2: 5868D030-0B97-43A1-BF70-B6A7F4569DFA.jpeg
5868D030-0B97-43A1-BF70-B6A7F4569DFA.jpeg
  14294   Wed Nov 14 14:35:38 2018 SteveUpdateALARM emergency calling list for 40m Lab

It is posted at the 40m wiki with Gautam' help. Printed copies posted around doors also.

  14293   Tue Nov 13 21:53:19 2018 gautamUpdateCDSRFM errors

This problem resurfaced, which I noticed when I couldn't get the single arm locks going.

The fix was NOT restarting the c1rfm model, which just brought the misery of all vertex FEs crashing and the usual dance to get everything back.

Restarting the sender models (i.e. c1scx and c1scy) seems to have done the trick though.

Attachment 1: RFMerrors.png
RFMerrors.png
  14292   Tue Nov 13 18:09:24 2018 gautamUpdateLSCInvestigation of SRCL-->MICH coupling

Summary:

I've been looking into the cross-coupling from the SRCL loop control point to the Michelson error point.

[Attachment #1] - Swept sine measurement of transfer function from SRCL_OUT_DQ to MICH_IN1_DQ. Details below.

[Attachment #2] - Attempt to measure time variation of coupling from SRCL control point to MICH error point. Details below.

[Attachment #3] - Histogram of the data in Attachment #2.

[Attachment #4] - Spectrogram of the duration in which data in #2 and #3 were collected, to investigate the occurrance of fast glitches.

Hypothesis: (so that people can correct me where I'm wrong - 40m tests are on DRMI so "MICH" in this discussion would be "DARM" when considering the sites)

  • SRM motion creates noise in MICH.
  • The SRM motion may be naively decomposed into two contributions -
    • Category #1: "sensing noise induced" motion, which comes about because of the SRCL control loop moving the SRM due to shot noise (or any other sensing noise) of the SRCL PDH photodiode, and
    • Category #2: all other SRM motion.
  • We'd like to cancel the former contribution from DARM.
  • The idea is to measure the transfer function from SRCL control point to the MICH error point. Knowing this, we can design a filter so that the SRCL control signal is filtered and summed in at the MICH error point to null the SRCL coupling to MICH.
  • Caveats/questions:
    • Introducing this extra loop actually increases the coupling of the "all other" category of SRM motion to MICH. But the hypothesis is that the MICH noise at low frequencies, which is where this increased coupling is expected to matter, will be dominated by seismic/other noise contributions, and so we are not actually degrading the MICH sensitivity.
    • Knowing the nosie-budget for MICH and SRCL, can we AC couple the feedforward loop such that we are only doing stuff at frequencies where Category #1 is the dominant SRCL noise?

Measurement details and next steps:

Attachment #1

  • This measurement was done using DTT swept sine.
  • Plotted TF is from SRCL_OUT to MICH_IN, so the SRCL loop shape shouldn't matter.
  • I expect the pendulum TF of the SRM to describe this shape - I've overlaid a 1/f^2 shape, it's not quite a fit, and I think the phase profile is due to a delay, but I didn't fit this.
  • I had to average at each datapoint for ~10 seconds to get coherence >0.9.
  • The whole measurement takes a few minutes.

Attachments #2 and #3

  • With the DRMI locked, I drove a sine wave at 83.13 Hz at the SRCL error point using awggui.
  • I ramped up the amplitude till I could see this line with an SNR of ~10 in the MICH error signal.
  • Then I downloaded ~10mins of data, demodulated it digitally, and low-passed the mixer output.
  • I had to use a pretty low corner frequency (0.1 Hz, second order butterworth) on the LPF, as otherwise, the data was too noisy.
  • Even so, the observed variation seems too large - can the coupling really change by x100?
  • The scatter is huge - part of the problem is that there are numerous glitches while the DRMI is locked.
  • As discussed at the meeting today, I'll try another approach of doing multiple swept-sines and using Craig's TFplotter utility to see what scatter that yields.

Attachments #2

  • Spectrogram generated with 1 second time strides, for the duration in which the 83 Hz line was driven.
  • There are a couple of large fast glitches visible.
Attachment 1: TF_sweptSineMeas.pdf
TF_sweptSineMeas.pdf
Attachment 2: digitalDemod.pdf
digitalDemod.pdf
Attachment 3: digitalDemod_hist.pdf
digitalDemod_hist.pdf
Attachment 4: DRMI_LSCspectrogram.pdf
DRMI_LSCspectrogram.pdf
  14291   Tue Nov 13 16:15:01 2018 SteveUpdateVACrga scan pd81 at day 119

 

 

Attachment 1: pd81-d119.png
pd81-d119.png
Attachment 2: pd81-560Hz-d119.png
pd81-560Hz-d119.png
  14290   Mon Nov 12 13:53:20 2018 ranaUpdateIOOloss measurement: oscope vs CDS DAQ

sstop using the ssscope, and just put the ssssignal into the DAQ with sssssome whitening. You'll get 16 bitsśšß.

Quote:

I increased the resolution on the scope by selecting Average (512) mode. I was a bit confused by this, since Yuki was correct that I had only 4 digits recorded over ethernet, which made me think this was an i/o setting. However the sample acquisition setting was the only thing I could find on the tektronix scope or in its manual about improving vertical resolution. This didn't change the saved file, but I found the more extensive programming manual for the scope, which confirms that using average mode does increase the resolution... from 9 to 14 bits! I'm not even getting that many.

 

  14289   Sat Nov 10 17:40:00 2018 aaronUpdateIOOIMC problematic

Gautam was doing some DRMI locking, so I replaced the photodiode at the AS port to begin loss measurements again.

I increased the resolution on the scope by selecting Average (512) mode. I was a bit confused by this, since Yuki was correct that I had only 4 digits recorded over ethernet, which made me think this was an i/o setting. However the sample acquisition setting was the only thing I could find on the tektronix scope or in its manual about improving vertical resolution. This didn't change the saved file, but I found the more extensive programming manual for the scope, which confirms that using average mode does increase the resolution... from 9 to 14 bits! I'm not even getting that many.

There's another setting for DATa:WIDth, that is the number of bytes per data point transferred from the scope.

I tried using the *.25 scope instead, no better results. Changing the vertical resolution directly doesn't change this either. I've also tried changing most of the ethernet settings. I don't think it's something on the scripts side, because I'm using the same scripts that apparently generated the most recent of Johannes' and Yuki's files; I did look through for eg tds3014b.py, and didn't see the resolution explicitly set. Indeed, I get 7 bits of resolution as that function specifies, but most of them aren't filled by the scope. This makes me think the problem is on the scope settings.

  14288   Sat Nov 10 17:32:33 2018 gautamUpdateLSCNulling MICH->PRCL and MICH->SRCL

With the DRMI locked, I drove a line in MICH using the sensing matrix infrastructure. Then I looked at the error points of MICH, PRCL and SRCL. Initially, the sensing line oscillator output matrix for MICH was set to drive only the BS. Subsequently, I changed the --> PRM and --> SRM matrix elements until the line height in the PRCL and SRCL error signals was minimized (i.e. the change to PRCL and SRCL due to the BS moving, which is a geometric effect, is cancelled by applying the opposite actuation to the PRM/SRM respectively. Then I transferred these to the LSC output matrix (old numbers in brackets).

MICH--> PRM = -0.335 (-0.2655)

MICH--> SRM = -0.35 (+0.25)

I then measured the loop TFs - all 3 loops had UGFs around 100 Hz, coinciding with the peaks of the phase bubbles. I also ran some sensing lines and did a sensing matrix measurement, Attachment #1 - looks similar to what I have obtained in the past, although the relative angles between the DoFs makes no sense to me. I guess the AS55 demod phase can be tuned up a bit.

The demodulation was done offline - I mixed the time series of the actuator and sensor signals with a "local oscillator" cosine wave - but instead of using the entire 5 minute time series and low-passing the mixer output, I divvied up the data into 5 second chunks, windowed with a Tukey window, and have plotted the mean value of the resulting mixer output.

Unrelated to this work: I re-aligned the PMC on the PSL table, mostly in Pitch.

Attachment 1: sensMat_2018-11-10.pdf
sensMat_2018-11-10.pdf
  14287   Fri Nov 9 22:24:22 2018 JonOmnistructure Wiring of Vacuum Acromag Chassis Complete

Wiring of the power, Ethernet, and indicator lights for the vacuum Acromag chassis is complete. Even though this crate will only use +24V DC, I wired the +/-15V connector and indicator lights as well to conform to the LIGO standard. There was no wiring diagram available, so I had to reverse-engineer the wiring from the partially complete c1susaux crate. Attached is a diagram for future use. The crate is ready to begin software developing on Monday. 

Attachment 1: IMG_2987.jpg
IMG_2987.jpg
Attachment 2: IMG_2985.jpg
IMG_2985.jpg
Attachment 3: IMG_2986.jpg
IMG_2986.jpg
  14286   Fri Nov 9 15:00:56 2018 gautamUpdateIOONo IFO beam as TT1 UL hijacked for REFL55 check

This problem resurfaced. I'm doing the debugging.

6:30pm - "Solved" using the same procedure of stepping through the whitening gains with a small (10 DAC cts pk) signal applied. Simply stepping through the gains with input grounded doesn't seem to do the trick.

Attachment 1: REFL55_wht_chk.png
REFL55_wht_chk.png
  14285   Wed Nov 7 23:07:11 2018 gautamUpdateLSCDRMI locking recovered

I had some success today. I hope that the tweaks I made will allow working with the DRMI during the day as well, though it looks like the main limiting factor in lock duty cycle is angular stability of the PRC.

  • Since there has been some change in the light levels / in vacuum optical paths, I decided to be a bit more systematic.
  • Initial guess of locking gains / demod phases was what I had last year.
  • Then I misaligned SRM, and locked PRMI, for the sideband resonant in the PRC (but still no arm cavities, and using 1f Refl error signals).
  • Measured loop TFs, adjusted gains, re-enabled boosts.
  • Brought the SRM back into the picture. Decided to trigger SRCL loop on AS110I rather than the existing POP22I (because why should 2f1 signal buildup carry information about SRCL?). New settings saved to the configure script. Reduced MICH gain to account for the SRC cavity gain.
  • Re-measured loop TFs, re-adjusted gains. More analysis about the state of the loops tomorrow, but all loops have UGF ~100-120 Hz.
  • Ran some sensing lines - need to check my sensing matrix making script, and once I get the matrix elements, I can correct the error signal demod phasing as necessary.

[Attachment #1]: Repeatable and reliable DRMI locks tonight, stability is mainly limited by angular glitches - I'm not sure yet if these are due to a suspect Oplev servo on the PRM, or if they're because of the tip-tilt PR2/PR3/SR2/SR3.

[Attachment #2]: A pass at measuring the TF from SRCL error point to MICH error point via control noise re-injection. I was trying to measure down to 40 Hz, but lost the lock, and am calling it for the night.

[Attachment #3]: Coherence between PRM oplev error point and beam spot motion on POP QPD.

Note that the MICH actuation is not necessarily optimally de-coupled by actuating on the PRM and SRM yet (i.e. the latter two elements of the LSC output matrix are not precisely tuned yet).

What is the correct way to make feedforward filters for this application? Swept-sine transfer function measurement? Or drive broadband noise at the SRCL error point and then do time-domain Wiener filter construction using SRCL error as the witness and MICH error as the target? Or some other technique? Does this even count as "feedforward" since the sensor is not truly "outside" the loop?

Attachment 1: Screenshot_from_2018-11-07_23-05-58.png
Screenshot_from_2018-11-07_23-05-58.png
Attachment 2: SRCL2MICH_crosscpl.pdf
SRCL2MICH_crosscpl.pdf
Attachment 3: PRCangularCoh_rot.pdf
PRCangularCoh_rot.pdf
  14284   Wed Nov 7 19:42:01 2018 gautamUpdateGeneralIFO checkup and DRMI locking prep

Earlier today, I rebooted a few unresponsive VME crates (susaux, auxey).

The IMC has been unhappy for a couple of days - the glitches in the MC suspensions are more frequent. I reset the dark offsets, minimized MCREFL by hand, and then re-centered the beam on the MC2 Trans QPD. In this config, the IMC has been relatively stable today, although judging by the control room StripTool WFS control signal traces, the suspension glitches are still happening. Since we have to fix the attenuator issue anyways soon, we can do a touch-up on IMC WFS.

I removed the DC PD used for loss measurements. I found that the AS beam path was disturbed - there is a need to change the alignment, this just makes it more work to get back to IFO locking as I have to check alignment onto the AS55 and AS110 PDs.

Single arm locking worked with minimal effort - although the X arm dither alignment doesn't do the intended job of maximizing the transmission. Needs a checkup.

PRMI locking (carrier resonant) was also pretty easy. Stability of the lock is good, locks hold for ~20 minutes at a time and only broke because I was mucking around. However, when the carrier is resonant, I notice a smeared scatter pattern on the ITMX camera that I don't remember from before. I wonder if the FF idea can be tested in the simpler PRMI config.

After recovering these two simpler IFO configurations, I improved the cavity alignment by hand and with the ASS servos that work. Then I re-centered all the Oplev beams onto their respective QPDs and saved the alignment offsets. I briefly attemped DRMI locking, but had little success, I'm going to try a little later in the evening, so I'm leaving the IFO with the DRMI flashing about, LSC mode off.

ELOG V3.1.3-